# Reasoning on Time-Series for Financial Technical Analysis

### v1

@@ 3.3. Time-Series Forecasting
Next, the projected time tokens $\mathbf{X}_{\text{time}}$ and aligned text tokens $\mathbf{X}_{\text{text}}$ are passed through consecutive LLM transformer blocks. To guide the modality alignment, after each transformer block in the temporal and textual branches, the embeddings are passed through a projection layer (chen2020simple) and matched via a feature regularization loss. This allows the forecasting model to tap into the deep contextual knowledge of the LLM while ensuring that the textual representations are consistent with the temporal dynamics at each layer. Formally, given $\mathbf{F}^{n}_{\text{time}}$ and $\mathbf{F}^{n}_{\text{text}}$, which are the outputs of the $n^{\text{th}}$ transformer block in the temporal and textual branches respectively, we define the feature regularization loss as:
(5) $\mathcal{L}_{\text{feature}}=\sum_{n=1}^{N}\gamma^{(N-n)}\,\text{sim}\bigl(\phi^{n}_{\text{text}}(\mathbf{F}^{n}_{\text{text}}),\phi^{n}_{\text{time}}(\mathbf{F}^{n}_{\text{time}})\bigr),$
@@
At the end of the transformer blocks, the features are passed through a final dense layer to produce the temporal- and text-based outputs, $\hat{\mathbf{y}}_{\text{time}}$ and $\hat{\mathbf{y}}_{\text{text}}$. These are also matched via $L_{1}$ loss:
(6) $\mathcal{L}_{\text{output}}=\text{sim}(\hat{\mathbf{y}}_{\text{time}},\hat{\mathbf{y}}_{\text{text}}).$
@@
3.4. Joint Conditional Training
@@
Figure 2. Joint conditional training component.
For this step, we first prompt for the reasoning output $\mathbf{o}$ using the previously fine-tuned time-series reasoning policy $\pi_{\theta}$. Next, we extract descriptive attribute classes $\mathbf{c}$ (i.e., its maximum, minimum, mean values) from the generated time-series $\hat{\mathbf{y}}_{\theta}$, which are analogous to the conditional class labels in text-conditional image generation (dhariwal2021diffusion). These are then used to condition the time-series forecasts via joint conditional training (see Figure 2): For each label, we concatenate it with the time-series forecasts $\hat{\mathbf{y}}_{\phi}(\mathbf{X})$ from the time-series forecasting model. These individual inputs pass through separate linear layers for fine-tuning, and are then aggregated via a projection layer to generate the conditioned time-series forecasts, which we denote here as $\hat{\mathbf{y}}_{\psi}(\mathbf{X},\mathbf{c})$.
Finally, we use a single neural network to parameterize both the conditional and unconditional models (ho2022classifier). Both the unconditional and conditional pipelines are trained concurrently by randomly setting $\mathbf{c}$ to the unconditional class identifier $\mathbf{\varnothing}$ with a predefined probability $p_{\text{uncond}}$, set as a hyperparameter. The overall model is then fine-tuned using MSE loss with the ground truth series $\mathbf{y}$:
(7) $\mathcal{L}_{\text{forecast}}(\phi)=\mathbb{E}_{\mathbf{X},\mathbf{y},\mathbf{c}}\left[\left\|\hat{\mathbf{y}}_{\psi}(\mathbf{X},\tilde{\mathbf{c}})-\mathbf{y}\right\|^{2}\right],$
(8) $\tilde{\mathbf{c}}\sim\begin{cases}\mathbf{c},&\text{with probability }1-p_{\text{uncond}}\\ \mathbf{\varnothing},&\text{with probability }p_{\text{uncond}}\end{cases}.$
@@
(9) $\hat{\mathbf{y}}=\hat{\mathbf{y}}_{\phi}(\mathbf{X})+s\cdot\bigl(\hat{\mathbf{y}}_{\psi}(\mathbf{X},\mathbf{c})-\hat{\mathbf{y}}_{\phi}(\mathbf{X})\bigr),$
@@
4. Experiments
Dataset: We evaluate our Verbal Technical Analysis (VTA) model extensively across multiple datasets. The first is the ACL18 StockNet dataset (xu2018stock), which includes historical price data for 88 U.S. stocks that are selected to represent the top 8-10 companies by market capitalization in each of the major industries. The dataset spans the period of 01/09/2012 to 01/09/2017. This dataset is a standard stock prediction benchmark that has been evaluated in multiple works (koa2023diffusion; feng2018enhancing; sawhney2020deep; feng2021time; li2023pen; chen2025dhmoe).
@@ 4. Experiments
Baselines: We compare against 12 state-of-the-art time-series methods: Transformer (vaswani2017attention), Reformer (kitaev2020reformer), Informer (zhou2021informer), Autoformer (wu2021autoformer), DLinear (zeng2023transformers), FiLM (zhou2022film), Crossformer (zhang2023crossformer), MICN (wang2023micn), LightTS (campos2023lightts), TimesNet (wu2022timesnet), TSMixer (chen2023tsmixer) and Non-Stationary Transformer (liu2022non). We also compare with two LLM-based time-series models: TimeLLM (jin2023time) and CALF (liu2025calf). These models are not explainable, including the two LLM-based models, which modify the embedding space for forecasting.
For evaluation against explainable models, we compare with reasoning LLMs: GPT-4.1 mini (openai2025gpt41) and DeepSeek-R1 (guo2025deepseek). To do so, we prompt these models to produce the time-series forecasts (gruver2023large; wang2024news) by reasoning on the time-series inputs.
Table 2. Performance comparison. The best baselines are underlined, and the best results are bolded.
@@
GPT-4.1 mini (openai2025gpt41) 0.0846 0.1827 0.4875 0.3191 0.0997 0.2128 0.1340 0.2358 0.2014 0.2376 0.4153 0.1072
DeepSeek-R1 (guo2025deepseek) 0.0788 0.1853 0.2920 0.3093 0.0776 0.2095 0.1227 0.2251 0.1428 0.2323 0.1750 0.0868
@@
Informer (zhou2021informer) 2.1846 0.9778 4.6823 1.1080 2.8986 1.0968 3.7031 1.1255 3.3672 1.0770 0.9650 0.8030
Transformer (vaswani2017attention) 1.5071 0.7762 3.9865 0.9784 2.1002 0.9042 2.8248 0.9194 2.6047 0.8946 0.9548 0.7628
Crossformer (zhang2023crossformer) 1.1848 0.6475 4.3396 0.9641 1.5656 0.7644 2.3503 0.8173 2.3601 0.7983 0.9501 0.7343
TSMixer (chen2023tsmixer) 1.4974 0.8193 3.5863 1.0905 1.1696 0.7227 2.5077 0.8473 2.1902 0.8700 0.9462 0.7561
Reformer (kitaev2020reformer) 1.1823 0.7628 2.4537 0.8503 1.5342 0.8302 2.4280 0.9048 1.8995 0.8370 0.9380 0.7465
LightTS (campos2023lightts) 0.6081 0.5213 1.0634 0.5509 0.6160 0.5128 1.0994 0.6103 0.8467 0.5488 0.8609 0.6134
DLinear (zeng2023transformers) 0.1589 0.3021 0.3880 0.4126 0.1716 0.3189 0.2407 0.3566 0.2398 0.3475 0.5088 0.3896
FiLM (zhou2022film) 0.0806 0.1927 0.2852 0.3085 0.0894 0.2157 0.1246 0.2370 0.1449 0.2385 0.1873 0.1103
Non-stationary (liu2022non) 0.0729 0.1822 0.2723 0.2993 0.0861 0.2079 0.1207 0.2305 0.1380 0.2300 0.1463 0.0776
MICN (wang2023micn) 0.0764 0.1878 0.2498 0.2922 0.0874 0.2108 0.1373 0.2405 0.1377 0.2328 0.1449 0.0888
Autoformer (wu2021autoformer) 0.0748 0.1866 0.2427 0.2947 0.0853 0.2103 0.1132 0.2273 0.1290 0.2297 0.0868 0.0765
TimesNet (wu2022timesnet) 0.0708 0.1789 0.2527 0.2902 0.0796 0.2022 0.1112 0.2203 0.1286 0.2229 0.0841 0.0482
@@
TimeLLM (jin2023time) 0.0704 0.1780 0.2439 0.2850 0.0776 0.1993 0.1127 0.2216 0.1262 0.2210 0.0663 0.0400
CALF (liu2025calf) 0.0674 0.1738 0.2412 0.2843 0.0762 0.1957 0.1092 0.2181 0.1235 0.2180 0.0460 0.0267
@@
Implementation Details: All LLMs used in VTA, including the reasoning model and transformer blocks for time-series forecasting, are trained using Low-Rank Adaptation (LoRA) (hu2022lora). Both input and output lengths $T$ and $T^{\prime}$ are set to 10, which is considered short-term forecasting in time-series works (li2022generative; liu2025calf). Technical Analysis is typically utilized for short-term stock trading (schwager1995technical).
For the reasoning model, we use Qwen2.5-7B-Instruct (qwen2.5) as our base model. For the forecasting model, we use GPT-2 (radford2019language) as the base model. For hyperparameters, we set the conditional probability $p_{\text{uncond}}$ to 0.3 and the guidance scale $s$ to 0.1.
5. Results
Performance Comparison. Table 2 reports the forecasting results.
@@ 5. Results
  • Among the time-series baselines, the biggest performance jump comes from the models which decomposes the data by trend and seasonality (i.e., FiLM, MICN, Autoformer and TimesNet). This could be attributed to the characteristics of stock prices, which are often affected by long-term and short-term business cycles (dalio2018principles). Another notable model that performs well is the non-stationary transformer, which might be attributed to the non-stationary behavior of stock price data (malkiel2019random).
@@
Figure 3. Correctness reward over steps.
Table 3. Ablation study of the LLM fine-tuning stages.
@@
Figure 4. An example of the reasoning trace and forecasts produced by our VTA framework.
@@
  • Finally, conditioning on the additional forecasting model helped to improve the performance further, showing the effectiveness of our combined verbal and latent VTA framework.
@@
Reasoning Quality. To evaluate the quality of the reasoning traces, we refer to past works on LLM explainability (koa2024learning; lin2024decoding) to design a set of relevant metrics for our task. Using these metrics, we surveyed 25 industry experts with professional experience in financial market analysis, with backgrounds from organizations such as JPMorgan, UBS, Evercore, and Allianz Global Investors, etc. For each, we presented the model outputs from VTA (ours), GPT-4.1 mini, and Deepseek-R1, showing both the forecasts and the textual reasoning. Respondents were blind to which model produced which output and were shown 15 randomly selected samples for evaluation. The samples were rated from 1 to 5.
Table 4. Comparison on reasoning quality.
Clarity Depth Accuracy Coherence Relevance
GPT-4.1 mini 3.43 1.88 2.15 3.15 2.53
Deepseek R1 3.07 2.20 2.36 3.03 2.56
VTA (Ours) 3.74 4.04 3.37 3.60 3.71
@@
From Figure 4, the results show that VTA achieved the highest average rating across all five metrics when compared to the other two LLMs. The best performance gains were observed in Depth, Accuracy, and Relevance. These criteria most directly reflect technical reasoning ability and the use of financial indicators. These findings suggest that our model, which was designed specifically for technical analysis, was able to produce reasoning outputs that were preferred by domain experts, demonstrating its practical strengths. The differences in Coherence and Clarity were smaller, which can be attributed to the fact that general-purpose LLMs like GPT-4.1 mini and Deepseek-R1 are good at producing fluent, well-structured text, even if they lack domain specialization. An example of the VTA reasoning outputs can be found in Figure 4.
@@
Portfolio Optimization. To justify the practical capability of VTA, we also evaluate the model in a real-life investment setting. Given that the model produces multi-step predictions, we are able to form portfolios that maximize the returns and minimize the volatility across the prediction length (koa2023diffusion). This is achieved by performing Markowitz optimization (selection1952harry) across the 10-day predictions. The portfolio is rebalanced daily, using the predicted returns and their covariance matrix.
@@
For evaluation, we compare against similar portfolios formed using the top-5 performing time-series models and all LLM baselines. The portfolio are compared on common investment metrics, such as their returns, volatility, maximum drawdown and Sharpe ratio (sharpe1994sharpe), which is a measure of risk-adjusted returns. We evaluate the portfolios across all 4 datasets, and report the average metrics.
From Table 5, we see that the portfolio constructed using VTA predictions demonstrates strong overall performance. It ranks as the strongest baseline on returns, volatility and maximum drawdown, and the values are very close to the top-performing models for each. Notably, VTA achieves the highest Sharpe ratio among all models, which represents the risk-adjusted returns. Given that the Sharpe ratio is one of the most common measure of investment performance, this highlights the practical effectiveness of VTA.
Table 5. Comparison on common portfolio metrics.
@@
6. Conclusion

### v2

@@ 3.3 Time-Series Forecasting
Next, the projected time tokens $\mathbf{X}_{\text{time}}$ and aligned text tokens $\mathbf{X}_{\text{text}}$ are passed through consecutive LLM transformer blocks. To guide modality alignment, after each transformer block in the temporal and text branches, the embeddings pass through a projection layer (Chen et al., 2020) and are matched via a feature regularization loss. This ensures that the text representations are consistent with the temporal dynamics at each layer. Formally, given $\mathbf{F}^{n}_{\text{time}}$ and $\mathbf{F}^{n}_{\text{text}}$, which are outputs of the $n^{\text{th}}$ transformer block in the temporal and text branches, we define the feature regularization loss as:
$\mathcal{L}_{\text{feature}}=\sum_{n=1}^{N}\gamma^{(N-n)}\,\text{sim}\bigl(\phi^{n}_{\text{text}}(\mathbf{F}^{n}_{\text{text}}),\phi^{n}_{\text{time}}(\mathbf{F}^{n}_{\text{time}})\bigr),$ (6)
@@
At the end of the transformer blocks, the features are passed through a final dense layer to produce the temporal-based and text-based outputs, $\hat{\mathbf{y}}_{\text{time}}$ and $\hat{\mathbf{y}}_{\text{text}}$. These are also matched via $L_{1}$ loss:
$\mathcal{L}_{\text{output}}=\text{sim}(\hat{\mathbf{y}}_{\text{time}},\hat{\mathbf{y}}_{\text{text}}).$ (7)
@@
3.4 Joint Conditional Training
Figure 2: Joint conditional training component.
@@ 3.4 Joint Conditional Training
For this step, we first prompt for the reasoning output $\mathbf{o}$ using the time-series reasoning policy $\pi_{\theta}$. Next, we extract descriptive attribute classes $\mathbf{c}$ (i.e., its maximum, minimum, mean values) from the generated time-series $\hat{\mathbf{y}}_{\theta}$, which are used to condition (Dhariwal and Nichol, 2021) the time-series forecasts via joint conditional training (see Figure 2): For each label, we concatenate it with the time-series forecasts $\hat{\mathbf{y}}_{\phi}(\mathbf{X})$ from the time-series forecasting model. These inputs pass through separate linear layers for fine-tuning, and are then aggregated via a projection layer to generate the conditioned time-series forecasts, which we denote as $\hat{\mathbf{y}}_{\psi}(\mathbf{X},\mathbf{c})$.
Finally, we use a single neural network to parameterize both the conditional and unconditional models (Ho and Salimans, 2022). Both the unconditional and conditional pipelines are trained concurrently by randomly setting $\mathbf{c}$ to the unconditional class identifier $\mathbf{\varnothing}$ with a predefined probability $p_{\text{uncond}}$. The model is then fine-tuned using MSE loss with the ground truth series $\mathbf{y}$. We have:
$\mathcal{L}_{\text{forecast}}(\phi)=\mathbb{E}_{\mathbf{X},\mathbf{y},\mathbf{c}}\left[\left\|\hat{\mathbf{y}}_{\psi}(\mathbf{X},\tilde{\mathbf{c}})-\mathbf{y}\right\|^{2}\right],$ (8)
$\tilde{\mathbf{c}}\sim\begin{cases}\mathbf{c},&\text{with probability }1-p_{\text{uncond}}\\ \mathbf{\varnothing},&\text{with probability }p_{\text{uncond}}\end{cases}.$ (9)
@@
$\hat{\mathbf{y}}=s\cdot\hat{\mathbf{y}}_{\psi}(\mathbf{X},\mathbf{c})+(1-s)\cdot\hat{\mathbf{y}}_{\theta}(\mathbf{X}),$ (10)
@@
4 Experiments
Dataset: We evaluate our Verbal Technical Analysis (VTA) model extensively across multiple datasets. The first is the ACL18 StockNet dataset (Xu and Cohen, 2018), which includes historical price data for 88 U.S. stocks that are selected to represent the top 8-10 companies by market capitalization in each of the major industries. The dataset spans the period of 01/09/2012 to 01/09/2017. This dataset is a standard stock prediction benchmark that has been evaluated in multiple works (Feng et al., 2018; Sawhney et al., 2020; Feng et al., 2021; Li et al., 2023; Koa et al., 2023; Chen and Wang, 2025).
@@ 4 Experiments
Baselines: We compare against 12 state-of-the-art time-series methods: Transformer (Vaswani et al., 2017), Reformer (Kitaev et al., 2020), Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), DLinear (Zeng et al., 2023), FiLM (Zhou et al., 2022), Crossformer (Zhang and Yan, 2023), MICN (Wang et al., 2023), LightTS (Campos et al., 2023), TimesNet (Wu et al., 2022), TSMixer (Chen et al., 2023) and Non-Stationary Transformer (Liu et al., 2022). We also compare with two LLM-based time-series models: TimeLLM (Jin et al., 2024) and CALF (Liu et al., 2025a). These models are not explainable, including the two LLMs, which modify the embedding space for forecasting.
For evaluation against explainable models, we compare with reasoning LLMs: GPT-4.1 mini (OpenAI, 2025) and DeepSeek-R1 (Guo et al., 2025). To do so, we prompt these models to produce the time-series forecasts (Gruver et al., 2023; Wang et al., 2024) by reasoning on the time-series inputs.
Implementation Details: All LLMs used in the VTA model, including the reasoning model and the transformer blocks for time-series forecasting, are trained using Low-Rank Adaptation (LoRA) (Hu et al., 2022). Both input and output lengths $T$ and $T^{\prime}$ are set to 10, which is considered short-term forecasting in time-series works (Li et al., 2022; Liu et al., 2025a). Technical Analysis is typically utilized for short-term stock trading (Schwager, 1995).
For the reasoning model, we use Qwen2.5-7B-Instruct (Team, 2024) as our base model. For the forecasting model, we use GPT-2 (Radford et al., 2019) as the base model. For hyperparameters, we set the conditional probability $p_{\text{uncond}}$ to 0.3 and the guidance scale $s$ to 0.1. More details on the experimental settings and computational resources used can be found in Appendix A.
5 Results
Table 2: Performance comparison. The best baselines are underlined, and the best results are bolded.
@@ 5 Results
GPT-4.1 mini (OpenAI, 2025) 0.0846 0.1827 0.4875 0.3191 0.0997 0.2128 0.1340 0.2358 0.2014 0.2376 0.4153 0.1072
DeepSeek-R1 (Guo et al., 2025) 0.0788 0.1853 0.2920 0.3093 0.0776 0.2095 0.1227 0.2251 0.1428 0.2323 0.1750 0.0868
@@
Informer (Zhou et al., 2021) 2.1846 0.9778 4.6823 1.1080 2.8986 1.0968 3.7031 1.1255 3.3672 1.0770 0.9650 0.8030
Transformer (Vaswani et al., 2017) 1.5071 0.7762 3.9865 0.9784 2.1002 0.9042 2.8248 0.9194 2.6047 0.8946 0.9548 0.7628
Crossformer (Zhang and Yan, 2023) 1.1848 0.6475 4.3396 0.9641 1.5656 0.7644 2.3503 0.8173 2.3601 0.7983 0.9501 0.7343
TSMixer (Chen et al., 2023) 1.4974 0.8193 3.5863 1.0905 1.1696 0.7227 2.5077 0.8473 2.1902 0.8700 0.9462 0.7561
Reformer (Kitaev et al., 2020) 1.1823 0.7628 2.4537 0.8503 1.5342 0.8302 2.4280 0.9048 1.8995 0.8370 0.9380 0.7465
LightTS (Campos et al., 2023) 0.6081 0.5213 1.0634 0.5509 0.6160 0.5128 1.0994 0.6103 0.8467 0.5488 0.8609 0.6134
DLinear (Zeng et al., 2023) 0.1589 0.3021 0.3880 0.4126 0.1716 0.3189 0.2407 0.3566 0.2398 0.3475 0.5088 0.3896
FiLM (Zhou et al., 2022) 0.0806 0.1927 0.2852 0.3085 0.0894 0.2157 0.1246 0.2370 0.1449 0.2385 0.1873 0.1103
Non-stationary (Liu et al., 2022) 0.0729 0.1822 0.2723 0.2993 0.0861 0.2079 0.1207 0.2305 0.1380 0.2300 0.1463 0.0776
MICN (Wang et al., 2023) 0.0764 0.1878 0.2498 0.2922 0.0874 0.2108 0.1373 0.2405 0.1377 0.2328 0.1449 0.0888
Autoformer (Wu et al., 2021) 0.0748 0.1866 0.2427 0.2947 0.0853 0.2103 0.1132 0.2273 0.1290 0.2297 0.0868 0.0765
TimesNet (Wu et al., 2022) 0.0708 0.1789 0.2527 0.2902 0.0796 0.2022 0.1112 0.2203 0.1286 0.2229 0.0841 0.0482
@@
TimeLLM (Jin et al., 2024) 0.0704 0.1780 0.2439 0.2850 0.0776 0.1993 0.1127 0.2216 0.1262 0.2210 0.0663 0.0400
CALF (Liu et al., 2025a) 0.0674 0.1738 0.2412 0.2843 0.0762 0.1957 0.1092 0.2181 0.1235 0.2180 0.0460 0.0267
@@
Performance Comparison. Table 2 reports the forecasting results. We can observe the following:
@@
  • Among the time-series baselines, the biggest performance jump comes from the models which decomposes the data by trend and seasonality (i.e., FiLM, MICN, Autoformer and TimesNet). This could be attributed to the characteristics of stock prices, which are often affected by long-term and short-term business cycles (Dalio, 2018). Another notable model that performs well is the non-stationary transformer, which might be attributed to the non-stationary behavior of stock price data (Malkiel, 2019).
@@
More details of the statistical significance of the experiments can be found in Appendix C.
Figure 3: Correctness reward over steps.
Table 3: Ablation study of the LLM fine-tuning stages.
@@
  • Finally, conditioning on the additional forecasting model helped to improve the performance further, showing the benefit of enhancing external verbal reasoning with internal latent understanding.
Contribution of Reasoning Component. To study the contribution of the reasoning component on forecasting performance, we artificially corrupt the reasoning trace to observe their impact on the final forecasts. This was done in two ways: (1) Adversarial: We add adversarial noise to the technical indicator values; (2) Remove Information: Within the prompt, we force the LLM to not utilize certain indicators in its reasoning. The forecasting performances are reported in Figure 4.
Figure 4: Change in reasoning performance when the reasoning traces are corrupted.
In Figure 4, removing the indicators leads to a clear degradation in performance, demonstrating that the reasoning traces provide genuinely useful guidance to the model. Adding adversarial noise also reduces overall performance but does not yield a consistent trend. A possible explanation is that, during joint conditional training, the model gradually learns to rely more heavily on the time-series forecasting component once it detects that the reasoning signals have become unreliable.
Reasoning Quality. To evaluate the quality of the reasoning traces, we refer to past works on LLM explainability (Koa et al., 2024; Lin et al., 2024) to design a set of relevant metrics for our task. Each reasoning trace and its associated time-series forecast was scored on a scale from 1 (poor) to 5 (excellent). The criteria are defined as follows:
@@
Using these metrics, we surveyed 25 industry experts with professional experience in financial market analysis, with backgrounds from organizations such as JPMorgan, UBS, Evercore, and Allianz Global Investors, etc. For each, we presented the model outputs from VTA (ours), GPT-4.1 mini, and Deepseek-R1, showing both the forecasts and the textual reasoning. Respondents were blind to which model produced which output and were shown 15 randomly selected samples for evaluation.
Figure 5: Performance of VTA on the financial time-series reasoning task.
@@
From Figure 5, the results show that VTA achieved the highest average rating across all five metrics when compared to the other two LLMs. The best performance gains were observed in Depth, Accuracy, and Relevance. These criteria most directly reflect technical reasoning ability and the use of financial indicators. These findings suggest that our model, which was designed specifically for technical analysis, was able to produce reasoning outputs that were preferred by domain experts, demonstrating its practical strengths. The differences in Coherence and Clarity were smaller, which can be attributed to the fact that general-purpose LLMs like GPT-4.1 mini and Deepseek-R1 are good at producing fluent, well-structured text, even if they lack domain specialization. More details on this, including statistical significance and open-ended responses, are found in Appendix D.1.
Case Studies. To illustrate the capabilities of VTA, we present some qualitative case studies to demonstrate its reasoning process. For the figures, the horizontal axis shows the time-steps while the vertical axis shows the scaled prices. In Figure 7, we see that VTA was able to correctly reason about the (downward) price correction and subsequent upward trend. In Figure 7, VTA correctly reason about the oscillating prices with a slight uptrend. More case studies can be found in Appendix E.
Figure 6: An example of VTA reasoning about a slight correction and possible rebound in price.
Figure 7: An example of VTA reasoning about prices oscillating within a range with a slight uptrend.
Generalization to Other Domains. In general, VTA is able to produce reasoning traces for any time-series data. To verify this, we run VTA on datasets from other domains. These include Healthcare (Wu et al., 2021) and Energy (Zhou et al., 2021), which contains time-series data on ILI (influenza-like illness) cases and oil temperature respectively. The generated reasoning traces are shown in Figure 8. In these domains, we observe that the time-series do not contain any intrinsic interpretable signals to reason over, and VTA do not go beyond simple trend extrapolation in its reasoning. It has also been shown in previous studies that large complex models do not meaningfully improve performance in these use cases (Tan et al., 2024). For these cases, VTA are still able to produce reasoning traces, but these do not contribute significant additional signals for the forecasting model. The forecasting performance of VTA on these datasets can be found in Appendix G.
Figure 8: Examples of VTA reasoning traces on the Healthcare and Energy domains.
@@
Portfolio Optimization. To justify the practical capability of VTA, we also evaluate the model in a real-life investment setting. Given that the model produces multi-step predictions, we are able to form portfolios that maximize the returns and minimize the volatility across the prediction length (Koa et al., 2023). This is achieved by performing Markowitz optimization (Markowitz, 1952) across the 10-day predictions. The portfolio is rebalanced daily, using the predicted returns and their covariance matrix.
@@
For evaluation, we compare against similar portfolios formed using the top-5 performing time-series models and all LLM baselines. The portfolio are compared on common investment metrics, such as their Sharpe ratio (Sharpe, 1994). The explanations of the portfolio metrics used are as follows:
  • Returns: Measure the percentage change in portfolio value over a given period, indicating overall profitability.
  • Volatility: Captures the dispersion of returns over time, with higher volatility reflecting greater fluctuations and uncertainty.
  • Maximum Drawdown: Represents the largest peak-to-trough decline in portfolio value, highlighting the worst observed loss during the evaluation window.
  • Sharpe Ratio: Assesses risk-adjusted performance by comparing excess returns to return volatility, where higher values indicate more efficient risk-taking.
We evaluate the portfolios across all 4 datasets, and report the average results across each metric.
Table 4: Comparison on common portfolio metrics.
@@
From Table 4, we see that the portfolio constructed using VTA predictions demonstrates strong overall performance. It ranks as the strongest baseline on the returns, volatility and maximum drawdown metrics, and the values are very close to the top-performing models for each. Notably, VTA achieves the highest Sharpe ratio among all models, which represents the risk-adjusted returns. Given that the Sharpe ratio is one of the most common measure of investment performance, this highlights the practical effectiveness of the VTA forecasts. More advanced financial analysis on the portfolios can be found in Appendix H.
6 Conclusion